OpenStudy (anonymous):

Find the cube roots of 27(cos 330° + i sin 330°).

OpenStudy (rational):

\[27^{1/3} \left[\cos\left(\frac{330+360k}{3}\right) +i\sin\left(\frac{330+360k}{3}\right) \right]\] plugin \(k=0,1,2\) and simplify

OpenStudy (anonymous):

z0=\[z0=\sqrt[3]{27}(\cos(\left[\begin{matrix}330 + 2(0)\pi \\ 3 & \end{matrix}\right] + i \sin \left[\begin{matrix}330 + 2(0)\pi \\ 3 & \end{matrix}\right])\]

OpenStudy (anonymous):

right ?